The table below lists the prices and quantises consumed of three different goods from 2010-2012. Good yPrice()Quantity Price (3)uantity 10 15 a For 2010, 2011, and 2012, determine the amount that a ypical consumer pays each year to purchate the qansted in the table abowe Instructions: Round your answers to the nearest whole number 2010 Consumer expenditure instructions: Round your answers to one decimal place b. The percentage change in the amount the consumer paid was . It is problematic to use your answers to part b as a measure of inflation because (both price and consumotion are cthanging d. Suppose we take 2010 as the base year, implying that the market basket is fxed at 2010 consumption levels. Using 2010 consumption levels, the rate of infation was 200 S 220 260 10 %from 2010 to 2011,and 136 %from 2011to2012. % from 2010 to 2011, and % from 2011 to 2012. Repeat the exercise from part d, now assuming that the base year is 2011. Using 2011 consumption levels, the rate of inflation from was 2011, and %from 2010 to % from 2011 to2012. f. Your answers from parts d and e were diferent because (Click to select)
Answer #1

a. Consumer Expenditure:

2010 : 10 × 10+5 × 18 1 × 10= 200

2011 : 15 × 8 3 × 30 2 5 220

2012 : 20 × 5+4 × 25+5 × 10 250

b. Percentage change in the amount the consumer paid from 2010 to 2011: 10%

220-200 200一一× 100

Percentage change in the amount the consumer paid from 2011 to 2012: 13.6%

250 220 220 --× 100

c. It is problematic to use your answers to part b as measure of inflation because both price and consumption are changing.

d. Using 2010 as base year,

2010 : 10 × 10+5 × 18 1 × 10= 200

2011 : 15 × 10+3 × 18+2 × 10= 224

2012 : 20 × 10+4 × 18 5 × 10 322

Rate of inflation from 2010 to 2011: 12%

224-200 × 100= 12 200

Rate of inflation from 2011 to 2012: 43.7%

322- 224 224 13 7 100 45

e. Using 2011 as base year,

2010 : 10 × 8 5 × 30 1 × 5 235

2011 : 15 × 8 3 × 30 2 5 220

2012 : 20 × 8+4 × 30 5 × 5 305

Rate of inflation from 2010 to 2011:-6.4%

rac{220-235}{235} imes100=-6.38

Rate of inflation from 2011 to 2012: 38.6

305 -220 220 --× 100:

f. Answers in part d and e were different because base years put different weights on the goods.